OpenStudy (sonyalee77):

Suppose you want to show the data in a circle graph. Find the measure of the central angle for 7th grade. Round your answer to the nearest degree, if necessary. Grade Number of Students 4th 45 5th 25 6th 60 7th 50 8th 60 9th 45

jimthompson5910 (jim_thompson5910):

how many students are there total

OpenStudy (sonyalee77):

285

jimthompson5910 (jim_thompson5910):

how many in 7th grade

OpenStudy (sonyalee77):

50

jimthompson5910 (jim_thompson5910):

so the percentage of 7th graders is (50/285)*100 = 17.54%

jimthompson5910 (jim_thompson5910):

Now take 17.54% of 360 to get 17.54% of 360 = 0.1754*360 = 63.144 So the central angle for the 7th graders is roughly 63.144 degrees

OpenStudy (sonyalee77):

thanx

jimthompson5910 (jim_thompson5910):

np